OpenStudy (anonymous):

Express the radical in simplest form: square root of 1/11?

OpenStudy (anonymous):

√(1/11)

OpenStudy (anonymous):

\[\sqrt{1\over11} = \sqrt{1\over11} \times \sqrt{11\over11} = {\sqrt{11}\over 11}\]

OpenStudy (anonymous):

thanks bro u saved my life ...

OpenStudy (anonymous):

No problem.